The market price of a stock is $21.22 and it just paid a dividend of $1.54. The required rate of return is 11.19%. What is the expected growth rate of the dividend?
Answer #1

This question requires application of constant growth dividend discount model according to which: Po - Divi T-9 Po = Price of

Div1 = Div0 * (1 + g)

21.22 - 1.54 * (1+0) 0.1119 – 9

2.3745 - 21.22g = 1.54 + 1.54g

0.8345 = 22.76g

g = 3.67%